History of the Princeton OSETI Project

July 1999 The thirty year old mirror was removed from the telescope early in the month and sent with the two undergrads Brian Bower and Melissa Kemp to H.L. Clausing Inc. in Skokie, Ill. for realuminizing. The two week adventure went smoothly, although we have yet to see the finished product since we are waiting for the floor to be refinished before reinstalling the mirror. Check out the pictures of the unloading process!
August 12, 1999 Initiation of Princeton's OSETI Web pages by Brian Bower '02.
November 1999 The new floor was completed early in November.
November 20, 1999 Mirror reinstalled today! Check out the installation pictures.
December, 1999 Used laser collimator to center secondary and adjust tilt of secondary.
January 5, 2000 First light with the resurfaced mirrors. Triangular images noted.
January 7, 2000 Relieve pinch on secondary, images greatly improved. (Bob Pascal's idea.)
January 11, 2000 Open House for Princeton OSETI Group
February 2000 Start refurbishment of basement control room. Dome painting started.
March 2000 Jarosik and Angelescu start computer telescope control project. Open House for Astrophysics and Physics departments.
April 2000 Testing CCD autoguider and large eyepieces obtained by Bob Cava. Good results.
May 2000 Norm Jarosik fixes "clunk" problem and tunes up telescope control program.
May 25 and 26, 2000 Open House for Alumni (reunions). Huge turnout.
June 2000 Work on autoguider, pointing program, dome control program, and focal reducer for CCD camera.
July 2000 Norm implements pointing correction model for telescope. (Corrects for flexure and mount misalignment. More tests on focal reducer designs. (Changes f/13.5 telescope to f2.2; plate scale from 0.16 arcsec/pixel to 1.0 arcsec/pixel.)
August 2000 Princeton alumnus gives $10,000 to Princeton to support the OSETI project!

Telescope pointing (open loop) is good to an accuracy of 15 arcsec.

Juan Burwell starts wiring the OSETI detector box.

Offset guider design is completed and drawings are sent to the physics machine shop.

Juan and Alex Willman take over the Webmaster duties and make significant updates to the Princeton OSETI Web site.

Someone hacks into the computer used to control the telescope; Norm and Alex repair the damage (all software-related) and secure the computer against break-ins.

Norm completes the first "production" version of the telescope control software.

September - October 2000 Norm J and Dan A continue to refine the computer-controlled telescope pointing system, correcting for telescope flexure and mount misalignment. The dead-reckoning accuracy is now ±15 arcsec.

Juan B performs a Hartman test on the optics. His results are surprisingly accurate (typical statistical errors of 0.01") and indicate some misalignment of optics. There is some spherical aberration, perhaps due to us not using the optimum telescope focus position. Coma is the main problem, perhaps due to a small misalignment of the secondary mirror. It was aligned with a laser, but not to the accuracy needed to correct the small coma that Juan measures. Fortunately, he finds a very small astigmatism, indicating that the mirrors don't have much potato-chip distortion.

The shop has finished with the offset guider that Dave W designed this summer.

Paul Horowitz' group has finished building and checking out our circuit board. As soon as we get our box wired, Chip Coldwell will bring it to Princeton for installation and final checks.

A New York Times (NJ Section) reporter wants to do a story on the OSETI project. She sent a photographer on short notice, and some team members were gathered at the last minute. (Dave apologizes that there wasn't time to notify everyone.)

Our SBIG ST-7E CCD camera has arrived. It seems to be working well. Norm says that nebulae and galaxies photograph quite well.

November 4 to 7 The offset guider is installed. Norm J and Alex W test the latest configuration and obtain several images, which are now online.
March 2001 Work on the OSETI detector has been slow throughout the winter, but significant progress is now being made toward completion.

We tried several times to have an Open House for the MacMillan folks who helped get the Observatory into shape, but the weather didn't cooperate with us, so we'll try again next month.

We had a Messier Marathon on the night of March 19-20. Click here to see what Bob Cava had to say about it.

We had an Open House for Friends of the 36" on March 23. Very few people showed up, but we had a fun evening anyway. We captured images of most objects that we saw, several of which are available online.

May 12 to 19 We cleaned up the observatory on May 12, in preparation for regular telescope operations. Dan Angelscu gave a demonstration of computer-controlled telescope operations to those who hadn't seen it yet. Training sessions for telescope operators who didn't participate in last year's sessions, should be held soon.

Chip Coldwell, a graduate student in Paul Horowitz' group at Harvard, came to Princeton and installed the circuit board in the Princeton OSETI detector. The box, power, and photon counting parts of the detector are finished. We get nice single photoelectron pulses and a tolerable amount of noise. We expect to install the OSETI detector very soon.

June 21 The Princeton OSETI detector is on the telescope! Dave and Norm spotted a few stars on Tuesday evening and aligned the OSETI aperture with the CCD camera, which will be used to center the target star on the OSETI detectors. The main unsolved problem at this point is how to keep the OSETI box dry to prevent arcing of the high voltage supply. We used most of a bottle of dry nitrogen Tuesday night. Hauling gas bottles to the observatory is a lot of work.

Norm and Dave have to concentrate on the MAP satellite (launch June 30) for a few weeks. After that we should start training sessions for group members interested in becoming OSETI operators (observers). We hope to be ready to go in September when the air is dry.

Norm and Chip Coldwell got the remote telescope control operating last week. The Princeton telescope can be set up to automatically track the Harvard telescope position when we wish to do simultaneous observations of a star list.

Bad news from our Harvard collaborators. Last week the Oakridge Observatory took the worst lightning strike on record. The telescope controls were knocked out, the OSETI box was damaged, and their GPS system probably has to be replaced.

We had Hugh Lippincott '03 working on the Princeton telescope for the past 3 weeks. He installed three TV cameras to monitor the telescope and shutter positions from the control room.

August 17 This hasn't been a good summer for observations, but Dave and Norm managed to get to the telescope a few times this month, and actually saw an OSETI-like pulse come in earlier this week! A few telescope and OSETI detector upgrades are still needed, but we are in a state where local observations can be made. We hope to start operator training sessions late next week.

Norm used the telescope to make movies of the Microwave Anisotropy Probe (the one that distracted us with a June 30 launch). He got MAP on its third "phasing loop", when it was almost out to the Moon's orbit. This week he got a faint, moving, blinking image even though the satellite is almost three lunar orbit radii away, on its way to L2.

August 24-26 The first set of operator training sessions this year occurred this weekend. More operator training sessions will be scheduled soon.

During these training sessions, we discovered a small optics misalignment and telescope imbalance. We plan on balancing the telescope again and creating a new telescope pointing model soon, to fix these problems.

Towards the end of each of the last two training sessions, Norm and Alex experimented with autoguiding the telescope, and got the autoguiding working well by the end of the last session.

If things continue to go as well as they have been, we expect to begin OSETI operations by the end of September.

August 30 Norm, Dave, and John Carlyle carefully balanced the telescope with the offset guider and OSETI box attached.
September 4-6, 14, 16-17 The second and third sets of operator training sessions went well. We now have at least 18 people qualified to operate the telescope and the OSETI detector.

Alex setup a new Web server for FitzRandolph Observatory and the Princeton OSETI project. The new Web pages are under construction and will be available soon.

October 5 We had an open house for the telescope operators and their guests. Only a few people showed up, but we had fun anyway.
October 13 We cleaned up the observatory, and made several minor (but useful) additions and adjustments, in preparation for regular operations.
October 18-31 Chip Coldwell, a Harvard graduate student, is implementing software to facilitate communications between the Princeton and Harvard OSETI computers, and is modifying our telescope control software to work properly with his software.

Norm and Chip updated our telescope pointing model, significantly improving our pointing accuracy. Norm continues to refine the pointing model.

Alex put the new OSETI and FitzRandolph Observatory web pages in production, on a new web server. The new OSETI web pages are accessible at the URL: http://observatory.princeton.edu/oseti

November 1 Alex and Norm installed a (relatively) new computer in the control room, from which we can control the CCD camera that is attached to our telescope. Chip and Norm got their software working well enough that we felt ready to try simultaneous OSETI observations with Harvard. We succeeded in obtaining a few sets of star coordinates from Harvard, pointing to the stars with very good accuracy, and monitoring the stars for OSETI-like pulses, all with minimal input from us. We plan on making a few more such preliminary observations soon.
November 17 Our first full night of simultaneous OSETI operations with Harvard.
2002 January 5 The weather has been better than we had expected for this winter, so we've been able to make 705 observations of 458 distinct stars over 221 hours 9 minutes. During those observations, we detected 2999 hits, 101 of which were "good" by Chip's criteria. Seven of our hits coincided (within three minutes) of Harvard's hits. One of these Harvard/Princeton coincidences were separated by only 16 seconds.

Norm has developed a Linux-based computer program that can work with his telescope control program to center a target star within our OSETI aperture and guide our telescope to keep the star centered, all with minimal input from a telescope operator. This program is being tested, and should be put into production soon. Norm's new guiding program may be enhanced to perform OSETI observations with practically no input from an operator. The telescope operators/observers would then have more time available for other pursuits, such as...

We now have the capability of making useful position and magnitude measurements of known asteroids and comets, and of searching for undiscovered asteroids and comets. Alex Willman and Joe Maffei have implemented the tools that we need for such a program of measurements and searches. Alex has made a few preliminary observations, and we expect our first set of observations to be submitted the International Astronomical Union's Minor Planet Center soon. Additional information about this new project is available on the Princeton Astrometry and Photometry Web pages.

February 18 Harvard has removed their OSETI detector from their telescope for testing of another instrument. Their OSETI detector will be off their telescope for about a week. There will be no coordinated OSETI observations with Harvard during this time.
February 24 Too many of our dome's retainer wheels have become stuck to allow us to rotate the dome safely. Ed and Norm will work on repairing these wheels. All observations must be put on hold until the retainer wheels have been repaired.
February 28 Ed, Norm, and the staff from the Physics Department Machine Shop have repaired six of the dome's retainer wheels. The dome rotation is now smoother than ever, so we can resume observations. The other fourteen retainer wheels should receive the same treatment too, so we'll take care of them this spring.
March 3 Harvard has put their OSETI detector back on their telescope, so we can resume coordinated OSETI observations with Harvard.
May 4 Chip Coldwell and Paul Horowitz discovered that their OSETI box is out of alignment. They suspect that it has been misaligned since June 2001. Paul and Chip have fixed this problem.

A small but vigorous crew cleaned up our dome again.

May 18 and 19 We held an open house for members of the Princeton University community on May 18, and another open house for a specific group on May 19. We plan on holding more open houses later this year.
June 15-18 A few of us have suspected for some time that the images we have been obtaining are not as good as they could be. Therefore, John Miller invited an optics expert, Dick Parker, to our observatory to examine our telescope and see if its performance could be improved. Dick discovered that our telescope's primary mirror was slightly misaligned towards the west, and its collimation needed to be adjusted. Dick also found that some light was reflecting from the telescope mirrors more than once, because the central hole within the primary mirror was smaller than the diameter of the secondary mirror. With Dick's guidance, an enthusiastic group of friends and observers finely collimated our primary mirror. Dave and Norm also added a black aluminum ring to the bottom of the baffle inside the telescope, blocking the light that would have reflected too many times. We have taken a few new images to test our telescope's new performance, and will examine the new images and put them online soon.
June 21-24 We had problems with the declination drive again. Norm repaired its motor and fine drive clutch, and it appears to be working again.

We also discovered that we did not reassemble the instrument rotator properly after the recent collimation, as we cannot lock it into position. We'll have to remove the rotator again, fix it, and reassemble it properly. Until then, Norm has rigged the rotator so it won't move.

July 6 The primary mirror has moved since the recent collimation. We apparently did something during the collimation, which has destabilized the kinematic mount on which the mirror sets. We can still perform OSETI observations, but this problem should be addressed. We may have to remove the instrument assembly, rotator, mirror and cell, service the mount, then reassemble everything.
July 11 Patrick Regan, senior correspondent for science and technology at NJN News, interviewed Dave, Alex, John, and Joe about the OSETI project.
July 12 Several observers met to discuss the state of the OSETI project and things that we wanted done. We made a list of tasks to be done and minor purchases, which should improve our operations. We also decided to monitor the effect that our moving primary mirror is having on our observations through July, then schedule maintenance on the kinematic mount during August.
July 29 - September 11 Severe storms and electrical power problems damaged part of our OSETI box, some of the electronics that control our observatory dome, and the computer that controls our telescope. We repaired and upgraded our electrical power infrastructure and the damaged components. During these repairs, we also took the opportunity to repair the instrument rotator and mirror support problems that we discovered earlier.
2002 September 5 David T. Wilkinson, leader of the Princeton OSETI project, died.
September 2002 through January 2003 The Facilities department built us a bathroom in the basement of the observatory. Now we don't have to visit another building when nature calls!
October 19 The observatory was opened to a group from the Jewish Center on October 19 from 8 to 10 PM.
November 2002 The November 2002 issue of Discover magazine contains an article by Michael Lemonick about OSETI and the Harvard-Princeton OSETI project.

We held an Open House at the observatory on November 9.

Chip Coldwell and his Harvard colleagues have put up webcams in their Oak Ridge Observatory so we can see whether they're performing OSETI observations.

December 11-16 A shorted synchro resolver caused some telescope pointing problems that prevented us from making OSETI observations. Norm has repaired the synchro.
2003 January 1 A scientific paper is being written on the Harvard-Princeton OSETI project.
January 11 Herman Ashley provided us with a color printer to use in our control room.
January 11-30 Exceptionally cold weather and accumulations of ice on our observatory dome put too much stress on our dome shutter mechanism, causing the shutter cables to come off their pulleys. This prevented us from performing OSETI observations. Norm and the Physics Department machine shop built "keepers" to hold the cables in place, and fixed the shutter mechanism.
February 16 Ed Bluvius, a wire rope engineer, examined the wire ropes and related mechanisms that move our observatory dome and floor. He found that some minor repairs should be sone soon, and other major work will need to be done sometime over the next few decades, but the wire ropes are in good shape overall.
February 22 The observatory was opened to a local boy scout troop.
March 15 The observatory was opened to a local cub scout troop on March 15 from 6:30 to 8:30 PM.
April 2003 The observatory was opened to visitors on April 12 from 8 to 10 PM.

The observatory was opened to Prof. Paul Steinhardt's Physics 111 class on the evening of April 30.

May 2003 The May 2003 issue of Discover magazine contains a letter from Eunice Wilkinson about the article in the November 2003 issue.

The observatory was opened to Princeton University Alumni during Reunions, on the evenings of May 29 and 30.

May 31 - June 4 A lightning strike near the observatory damaged the dome controls on May 31. Norm replaced the damaged components on June 4.
July 25-29 The hard drive in the computer that controls our CCD camera stopped working on July 25. Norm replaced the hard drive, and Norm and Alex reinstalled the software needed on that computer. No OSETI observations took place during this time.
August 2003 Alex Willman replaced the computer that observers use to access the Internet, with a much faster one that also has a DVD-ROM drive.

The observatory was opened to invited visitors on August 27 and 28 from 9:30 to 11:30 PM to view Mars during its historically closest approach.

September 2003 Maggie Kirkland and James Wray are leading an effort to recruit new student volunteers.

The Department of Astrophysical Sciences' open house on September 7 was held at FitzRandolph Observatory instead of at Peyton Hall because of accidental damage to the equipment in the Peyton Hall Observatory.

The observatory lost power from the evening of September 18 through the evening of September 19, due to the effects of tropical storm (former hurricane) Isabel.

October 2 The observatory was opened to a local boy scout troop on October 2 from 7 to 10 PM.
November 2003 Alex Willman compiled a list of people who have contributed to the OSETI project in one way or another.

The Times of Trenton published another article about the Princeton OSETI project, written by Elizabeth Landau.

2004 January 30 The observatory was opened to visitors on January 30 from 7 to 10 PM.
February 2 The observatory was opened to a class from Stuart Day School, as part of Project ASTRO Nova.
March 2004 The New Jersey Section of the American Association of Physics Teachers visited the observatory on the evening of March 12.

A technical paper describing the Harvard/Princeton OSETI projects and summarizing the results through November 2003 was submitted to The Astrophysical Journal on March 16 for review and possible publication.

The observatory was opened to a local girl scout troop on March 28 from 7 to 9 PM. Unfortunately, the sky was too cloudy for them to see anything through the telescope.

Aaron discovered a minor problem with the telescope's declination pointing on March 12, after the NJAAPT left the observatory. Norm found a few broken components, and replaced them on March 31.

April 2004 A tour of the observatory was given to Woodrow Wilson School faculty, staff, and their children on the afternoon of April 22.

The observatory was opened to a third grade class from the Community Park School on April 22 from 8 to 9:30 PM.

The observatory was opened to visitors on April 30 from 8 to 10 PM.

June 2004 Andrew Howard has received several positive responses from reviewers to the technical paper that he submitted for publication. He is modifying the paper according to the reviewers' suggestions, and will resubmit a new version of the paper soon.

The technical paper was resubmitted on June 14 and accepted for publication in The Astrophysical Journal!

October 1 A technical paper describing the Harvard/Princeton OSETI projects and summarizing the results through November 2003 was published in The Astrophysical Journal!
May 2005 The observatory where the Harvard OSETI group performed their observations has been closed, so the targeted OSETI search performed by the Harvard and Princeton OSETI groups has stopped, likely for good. The Harvard OSETI group plans on continuing its all-sky OSETI survey.

The observatory was opened to the Lawrence High School Science Olympians on May 10 from 8:30 to 10:00 PM.

The observatory was opened for Reunions on May 26 and 27 from 8:00 to 10:00 PM.

June 30 WZBN News interviewed Ed Groth, Alex Willman, and John Miller regarding the Princeton OSETI project and FitzRandolph Observatory.
July 2005 The hard drive in the Internet-connected observatory computer failed, so Alex W replaced the computer.
August 2005 A lightning strike appears to have severely damaged the computer and much of the electronics that control the telescope. Norm Jarosik indicates that getting things working again will require a major commitment of time, so the computer control system is down indefinitely.

Princeton OSETI · What is OSETI? · Latest News · Technical Terms · FitzRandolph Observatory

Please direct comments about these Web pages to oseti@Princeton.EDU